|x1-x2|= |x1|-|x2| for all x1,x2€R.
Is the statement true or false? Justify your answer.
1
Expert's answer
2019-05-01T10:11:57-0400

For all "a,b\in R" we have the triangle inequality

"|a|+|b|\ge |a+b|"

Setting "a={{x}_{1}}-{{x}_{2}}" and "b={{x}_{2}}", we obtain

"|{{x}_{1}}-{{x}_{2}}|+|{{x}_{2}}|\ge|{{x}_{1}}-{{x}_{2}}+{{x}_{2}}|"

That is

"|{{x}_{1}}-{{x}_{2}}|\ge |{{x}_{1}}|-|{{x}_{2}}|"

It follows that there exist x1 and x2 for which inequality

"|{{x}_{1}}-{{x}_{2}}|>|{{x}_{1}}|-|{{x}_{2}}|"

holds. For example, if x1=2 and x2=-3 we have

"|2-\left( -3 \right)|>|2|-|-3|"

or

"|5|>|2|-|3|"

that is

"5>-1"

Hence, the statement

"|{{x}_{1}}-{{x}_{2}}|\,=\,|{{x}_{1}}|\,-|{{x}_{2}}|"

is false.